OpenStudy (anonymous):

theta = 2. It's a polar curve question, I need to graph the curve. Help!!! Thank you.

OpenStudy (anonymous):

I just need to know how to find the radius. Thank you!

OpenStudy (anonymous):

the equation is just \[\theta = 2\] ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

that's just a straight line with angle 2 radians

OpenStudy (anonymous):

Oh. I didn't thought of that! Thanks man!

OpenStudy (anonymous):

np. if you want to convert it to rectangular coordinates, think of the relationship tan(theta) = y/x yielding y = x tan(2) notice how this is just a straight line, also, in rectangular coordinates with slope tan(2)

OpenStudy (anonymous):

I don't get it. How is it a straight line. I do know the relationship but, I'm confused.

OpenStudy (anonymous):

if you're only given an angle measurement, there is no bound on the radius. it can be anything. if you were asked to graph "45 degrees", for example, it would just be a line.

OpenStudy (anonymous):

Ohhhhhhhhh. It all makes sense now.

OpenStudy (anonymous):

Thanks man.